Consider two countries that produce cloth and widgets with labour as the unique production factor using a linear technology. Given the following information: Unit Produced by One Worker/Hour Cloth Widgets Home 200 400 Foreign 120 60 i. What is the opportunity cost of cloth in terms of widgets for the Home country? For the Foreign country? (5 points) ii. In which good does the Home country have comparative advantage? Why? (5 points) ii. Assume that on the world market one unit of cloth trades for widgets at an equilibrium price of P/Pw-1. If the two countries trade the two goods with each other following the Ricardian model of comparative advantage, show that both countries will gain. (10 points) At the world equilibrium price (Po/Pw-1), what is the relative wage wH/w between the Home country and the Foreign country? (10 points) iv. Assuming that the equilibrium price Pe/Pw is not known, could you provide a lower and an upper bound for the relative wage WH/w*? Explain! (10 points) v. vi. Would it make any difference to anything whether we double the Home countrys maximum outputs of cloth and widgets by doubling Homes labour supply or by a. b. doubling its labour productivity in cloth and widgets production? Discuss your answers with regard to the trade pattern, and to the wage ratio (10
Answer #1

(i) Opportunity cost (OC) is defined as the amount of a commodity given up to consume another commodity.

For Home country, in terms of widgets OC is 400/200 = 2 i.e. 1 cloth is given up for 2 of widgets

(ii) Comparative advantage is the ability to produce a good more efficiently than others.

Home is (200/120=5/3) times better at cloth while (400/60=20/3) better at widget.

So,home has comparative advantage in widgets. Thus foreign has comparative advantage in cloth.

(iii) Following Ricardian model,

Home will produce and export Widgets while foreign will produce and export cloth.

cost ratio for widget : 400/6 = 20/3 cost ratio for cloth = 200/120= 5/3 as shown above.

In absence of trade in Home, 200C/400W = 0.5C :1 W. As TOT is 1:1 (given), trade causes Home to consume 0.5 units of C more. Similarly for foreign,when no trade, 60W/120C = 0.5 2 : 1 C. After trade foreign can consume 0.5 units of W more. So, trade beneficial for both.

When countries trade using Ricardian model, both the country's consumption frontier expands as it exports the commodity it has CA in and imports the other commodity more in quantity than it would do if it produced alone.

(iv) For cloth, WH/WF = 200/120 = 5/3 . For Widget, WH/WF = 400/60 = 20/3

vi)

a) No. As doubling both cloth and widget production simultaneously will not change the ratio of OC.

b) Same as (a).